Refrigerated food is kept between -2 degrees Celsius and 14 degrees Celsius. The number of bacteria that would be present in the food can be calculated using the equation: N = 20t2 - 20t + 120, where N = number of bacteria in food and t = temperature in degrees Celsius At what temperature will the number of bacteria be minimal?
Answer #1

The given equation is equation of upward parabola

since a=coefficient of t^2 which is 20 which is greater than 0.

The minimal of upward parabola is occurred at its vertex.

The formula to find vertex is (-b/ 2a, f(-b/2a)).

Here a=20, b=-20, c=120.

-b/2a=-(-20)/40=1/2.

Vertex=(1/2, f(1/2)).

Vertex =(0.5,115).

At t= 0.5 degree Celsius the bacteria is minimum.